OpenStudy (lin.ivory):

use the fact that f(x)= U(x)/V(x) can be written as f(x)V(x) = U(x) and the product rule for derivatives to verify the quotient rule for derivatives.

OpenStudy (anonymous):

f'(x)V(x)+V'(x)f(x)=U'(x) f'(x)V(x)=U'(x)-V(x)f(x) f'(x)V(x)=U'(x)-V'(x)(U(x)/V(x)) same denominator f'(x)V(x)=[U'(x)V(x)-V'(x)U(x)]/V(x) f'(x)=[U'(x)V(x)-V'(x)U(x)]/[V(x)V(x)] f'(x)=[U'(x)V(x)-V'(x)U(x)]/[V(x)]^2

OpenStudy (lin.ivory):

so basically you started with substitution and you just substituted U(x) in?

OpenStudy (anonymous):

you have f(x)V(x)=U(x) take derivative of both sides and solve for f'(x) derivative of the left side requires product rule so you do that... :)